# DreamForge: Motion-Aware Autoregressive Video Generation for Multiview Driving Scenes

### v3

@@ Abstract
footnotetext: Corresponding author
@@ A.4 More Results
Simulation within DriveArena. We illustrate the visualizations of the simulations within DriveArena. Our DreamForge receives sequential data from the Traffic Manager in DriveArena and sends generated scenes of keyframes to the driving agents at 2 Hz, as stated in the implementation details. Without loss of generality, we present cases where the driving agents UniAD [7] and VAD [8] operate in an open-loop mode in Fig. 12. The results indicate that predictions from the driving agents regarding the road network and vehicle tracking are fundamentally accurate, with UniAD demonstrating superior perception results on the generated scenes.

### v4

@@ Abstract
footnotetext: * Equal contribution  ✉ Corresponding author
@@ A.2.2 Training and Inference
Mask‐Shift Mechanism in DreamForge-DiT
We employ a hybrid masking approach at the feature level during training in DreamForge-DiT, randomly selecting one of two masking strategies for each training sample (50% probability for each).
  • Random Masking: Following Open-Sora 1.2 [28], a random subset of frame positions in the sequence are designated as targets (mask set to True), while the remaining frames serve as context (mask set to False). This strategy encourages the model to learn generation conditioned on diverse temporal contexts.
  • Autoregressive Masking: The first $N$ frames are fixed as context (setting them to False in the mask) ,and the model is tasked with generating the subsequent $T-N$ frames (setting them to True), enforcing sequential, autoregressive generation from past to future.
By alternating between these two strategies during training, we found that DreamForge-DiT demonstrates superior performance in long video generation, producing sequences that effectively maintain both local detail and global coherence across extended autoregressive temporal horizons.
@@ A.4 More Results
Scenario Base Model NC $\uparrow$ DAC $\uparrow$ EP $\uparrow$ TTC $\uparrow$ C $\uparrow$ PDMS $\uparrow$
DriveArena SD 1.5 0.792 0.942 0.738 0.771 0.749 0.636
DreamForge SD 1.5 0.843 0.958 0.728 0.829 0.704 0.669
DreamForge DiT, 3D VAE 0.8721 0.9220 0.7815 0.829 0.693 0.6753
Table 10: Open-loop evaluation results. The bold means the best result.
Scenario Method PDMS $\uparrow$ ADS $\uparrow$
singapore-
onenorth
DriveArena 0.7615 0.1282
DreamForge 0.8102 0.1197
DreamForge-DiT 0.8071 0.1207
boston-
seaport
DriveArena 0.4952 0.0450
DreamForge 0.7401 (+ 0.2449) 0.0760 (+ 0.0310)
DreamForge-DiT 0.8192 (+ 0.3240) 0.1448 (+ 0.0998)
Table 11: Close-loop evaluation results with UniAD in DriveArena
@@
Simulation within DriveArena. We illustrate the visualizations of the simulations within DriveArena. Our DreamForge receives sequential data from the Traffic Manager in DriveArena and sends generated scenes of keyframes to the driving agents at 2 Hz, as stated in the implementation details. Without loss of generality, we present cases where the driving agents UniAD [7] and VAD [8] operate in an open-loop mode in Fig. 12. The results indicate that predictions from the driving agents regarding the road network and vehicle tracking are fundamentally accurate, with UniAD demonstrating superior perception results on the generated scenes. Additionally, we integrated with the DriveArena [13] simulation platform to demonstrate the practical benefits of DreamForge-DiT. In both open-loop and closed-loop evaluations as shown in Table 10 and 11, the model provided more reliable and realistic scenarios for testing vision-based driving agents. The enhanced temporal coherence and controllability of DreamForge-DiT resulted in more accurate simulations of real-world driving conditions.